OpenStudy (anonymous):

A teacher gave his class two quizzes; 80% of the class passed the first quiz, but only 60% of the class passed both quizzes. What percent of those who passed the first quiz also passed the second quiz?

OpenStudy (anonymous):

is it 20%?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

thanks oops lol

OpenStudy (anonymous):

60/80 = 75%

OpenStudy (anonymous):

i think it might be 75 like @mangorox said

OpenStudy (anonymous):

Jean bought a bag of candy that contained red candies, green candies, and brown candies. Some candies are marked with a smile and others are not marked. The probabilities are displayed in the table below: Red Green Brown Smile 0.10 0.30 0.10 No smile 0.20 0.15 0.15

OpenStudy (anonymous):

If a random candy is chosen from this bag, what is the probability that the candy was red or had no smile?

OpenStudy (anonymous):

@mangorox @OOOPS

OpenStudy (anonymous):

whoops its 0.60 0.20 0.30 0.50

OpenStudy (anonymous):

To make a decimal into a percent, move the decimal dot two places to the right. So, to probability to choose a red candy would be: \[0.10 + 0.20 = 0.30 \approx 30 percent\] 30% red candy to be drawn out of the bag is the probaability

OpenStudy (anonymous):

okay i get it i think

OpenStudy (anonymous):

For candy to have no smile: \[0.20 + 0.15 + 0.15 = 0.50 \approx. 50percent\] 50% is the probability a candy with no smile will be drawn

OpenStudy (anonymous):

*50percent

OpenStudy (anonymous):

wait so what is the answer

OpenStudy (anonymous):

if its both 30% and 50%

OpenStudy (anonymous):

weird ill put 30